OpenStudy (anonymous):

PLEASE HELP!!! (Algebra, lesson Fractional Coefficients) y = ¼ x + 7 y = ½ x + 5 The y-coordinate of the solution to the system shown is multiple choice as 9, 11 ½ , or 19. I have tried multiple times to solve this problem. How do I solve it correctly?

OpenStudy (anonymous):

½ x + 5 = ¼ x + 7 ¼ x - ½ x = 5-7 -¼ x = -2 x= 8 so y=¼ (8)+7= 9

OpenStudy (johnweldon1993):

we just went through this problem @ScienceQs the answer would be 9 just like before